Freddie and Jason have just opened the Texas Toothpick, a chain-saw sharpening and repair service located...

  1. Freddie and Jason have just opened the Texas Toothpick, a chain-saw sharpening and repair service located on Elm Street. The Texas Toothpick promises same-week repair service. Freddie and Jason are concerned that a projected dramatic increase in demand as the end of October nears will cause service to deteriorate. Freddie and Jason have had difficulty attracting employees, so they are the only workers available to complete the work.

Safety considerations require that they work together for no more than 80 hours per week. The first chain-saw sharpening and repair required 10 hours of work, and an 80 percent learning curve is anticipated.

Week                             Units               Cumulative Units

October 2–6                     8                                8

October 9–13                 19                               27

October 16–20               10                               37

October 23–27               27                               64

  1. How many total hours are required to complete the 37th unit? (1pt)
  2. Will Freddie and Jason be able to keep their same week service promise during their busiest week just before Halloween? (2pts)
Answer #1
  1. How many total hours are required to complete the 37th unit?

Answer: 233.67 hours

Calculation

To calculate the total hours to complete the required units of 37, first we need to calculate each weeks required hours.

It is mentioned that first chain saw requires 10 hours which includes 80% of learning curve.

That is, 80% of the time srequired for learning and 20% required for performing the task.

First chain Saw sharpening and repair hours required = 10 hours

Then we need to find the  Learning Curve Rate from the Conversion Factors for the Cumulative Average Number of Direct Labor Hours per Unit.

So under the 80% learning curve, we need to look up the number of units in the n column and then find the rate. So the rates for units will be:

Week 1: 8 units = .66824

Week 2: 19 units = .53178

Week 3: 10 units = .61354

Week 4: 27 units = .48167

So then we need to multiply it with the  First chain Saw sharpening and repair hours required.

Week 1: 10 * .66824 = 6.6824

Week 2: 10 * .53178 = 5.3178

Week 3: 10 * .61354 = 6.1354

Week 4: 27 * .48167 = 4.8167

Then we need to multiply it with the number of units cumulated and then deduct with previous cumultive result

Week 1:6.6824 * 8 = 53.46

Week 2: 5.3178 * 27 = 143.58 - 53.46 = 90.12

Week 3: 6.1354 * 37 = 233.67 - 143.58= 90.09

Week 4: 4.8167 * 64 = 308.27 - 233.67 =74.60

So the total required for 64 units = 53.46 + 90.12 + 90.09 + 74.60 = 308.27

Here we are asked to find the hours required to complete the 37th unit, Here we need to sum up till the 3rd week. So,

= 53.46 + 90.12 + 90.09 = 233.67

2. Will Freddie and Jason be able to keep their same week service promise during their busiest week just before Halloween?

Yes,

Freddie and Jason has planned to work together for no more than 80 hours per week. It is due to safety considerations. So, the week before Halloween is the Week 4, where they are planning to produce highest number of units which is 27. As we have calculated above, the hours required is 74.60, which is less than the planned 80 hours. So Freddie and Jason will be able to keep their same week service promise during their busiest week just before Halloween. But for the weeks before the Week 4, that is week 2 and week 3 they wont be able to keep their same week service promise as the week 4 because the hours working will be 90 + hours.